I really don't understand this question:
The cost to rent a car is given by the formula C= 50d + 0.2k,
where d = the number of days rented
and
k= the number of km driven.
Lin has $300 to spend on car rental for her 4 day holiday. How far can she travel on this holiday?

Answer by longjonsilver(2297)

well, the cost is $300 (this is C), number of days rental is 4 (this is d). Find k.

C= 50d + 0.2k
300 = 50*4 + 0.2k
300 = 200 + 0.2k
100 = 0.2k
k = 100/0.2
k = 500km.

jon.